OpenStudy (anonymous):

The cost to produce a product is modeled by the function f(x) = 5x^2 − 70x + 258 where x is the number of products produced. Complete the square to determine the minimum cost of producing this product. A. 5(x − 7)^2 + 13; The minimum cost to produce the product is $13. B. 5(x − 7)^2 + 13; The minimum cost to produce the product is $7. C. 5(x − 7)^2 + 258; The minimum cost to produce the product is $7. D. 5(x − 7)^2 + 258; The minimum cost to produce the product is $258.

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

how far did you get?

OpenStudy (anonymous):

This was my first question like this

jimthompson5910 (jim_thompson5910):

what is the axis of symmetry in this case?

OpenStudy (anonymous):

5(x-7)^2

jimthompson5910 (jim_thompson5910):

hint: the axis of symmetry of ax^2 + bx + c is x = -b/(2a)

jimthompson5910 (jim_thompson5910):

what is the axis of symmetry of 5x^2 − 70x + 258

OpenStudy (anonymous):

5?

jimthompson5910 (jim_thompson5910):

what are 'a' and 'b' in this case?

OpenStudy (anonymous):

5 and 70

jimthompson5910 (jim_thompson5910):

close a = 5 b =-70

jimthompson5910 (jim_thompson5910):

now plug those into x = -b/(2a)

OpenStudy (anonymous):

Negative 700?

OpenStudy (anonymous):

i dont think thats right

jimthompson5910 (jim_thompson5910):

x = -b/(2a) x = -(-70)/(2*5) x = ???

OpenStudy (anonymous):

-70?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

-7?

jimthompson5910 (jim_thompson5910):

try again

OpenStudy (anonymous):

negative 70 divided by ten no?

OpenStudy (anonymous):

postive 7?

jimthompson5910 (jim_thompson5910):

yes it's positive 7

jimthompson5910 (jim_thompson5910):

so h = 7

jimthompson5910 (jim_thompson5910):

now that we know x = 7, plug it into the equation and tell me what you get

OpenStudy (anonymous):

608?

jimthompson5910 (jim_thompson5910):

that's way too high

OpenStudy (anonymous):

5(7)^2 − 70(7) + 258 <--- plug it in like that right?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

giving you what?

OpenStudy (anonymous):

is it anywhere in the hundreds?

OpenStudy (anonymous):

you said 608 is high

jimthompson5910 (jim_thompson5910):

yeah 608 is way too high

OpenStudy (anonymous):

so if it was 300 or 400 it'd still be too high?

jimthompson5910 (jim_thompson5910):

yes, it's a 2 digit number

OpenStudy (anonymous):

13?

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so that's why the equation turns into f(x) = 5(x-7)^2 + 13

OpenStudy (anonymous):

Thank you, thanks alot

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

thats wrong

OpenStudy (gahm8684):

@stabar whats the answer then